OpenStudy (anonymous):

two forces P and Q pass through a point A which is 4 ft to the right of and 3 ft above a moment center O. force P is 200 lb directed up to the right at 30degrees with the horizontal and force Q is 100 lb directed up to the left at 60degrees with the horizontal, determine the moment of the resultant of these two forces with respect to O.

OpenStudy (anonymous):

Drawing the situation described in the problem is extremely helpful. I've included a sketch of what things look like (bear with me, it's not a high quality image). We're given P & Q and asked to find their resultant; so we have to find their x and y components first. P = 200 lb Px = P * cos 30 = 200 * cos 30 Py = P * sin 30 = 200 * sin 30 Q = 100 lb Qx = Q * cos 60 = 100 * cos 60 Qy = Q * sin 60 = 100 * sin 60 It's asking for a resultant, Fr, of the two forces, so I'll make a table to help you see what's going on. Force | x component | y component P + 200 * cos 30 + 200 * sin 30 Q - 100 * cos 60 + 100 * sin 60 __________________________________________________________________________ Fr +200 * cos 30 - 100 * cos 30 +200 * sin 30 + 100 * sin 60 Note: Qx is negative because of my sign convention. I'm considering the positive x and positive y directions as positive. Qx is pointing in the negative x direction so we have to include a negative sign with Qx. Now you can calculate the x and y components of the resultant force, Fr. I'm not yet done but I want to make sure you're following along. If you understand what I just did, which direction is the x component of Fr pointing in and which direction is the y component of Fr pointing in?